(d^3-5d^2+7d-3)y=0

1
Expert's answer
2022-03-15T15:11:11-0400

The characteristic equation is

"D^3\u22125D^2+7D\u22123=0 \n\n\u27f9D^3\u2212D^2\u22124D^2+4D+3D\u22123=0\n\n\u27f9D^2(D\u22121)\u22124D(D\u22121)+3(D\u22121)=0"

"(D\u22121)(D^2\u22124D+3)=0 \n\n\u27f9(D\u22121)(D\u22121)(D\u22123)=0\n\n\u27f9D=1\u2228D=1\u2228D=3"


Therefore the general solution is

"y=c_1e^x+c_2xe^x+c_3e^{3x}=(c_1+xc_2)e^x+c_3e^{3x }"